I turned off the heater control for EY MN TILT OpLev for reducing glitches.
EY MN OpLev is so glitchy today (fig1), so we proably need to find the good temperature.
I made a script named search_good_OpLev_temp.py and stored it at /users/ushiba/script/python/.
This script changes the target temperature of temperature control every 0.1 degrees from 23.1 to 24.1.
At each temperature setting, script waits for 1 hour to stabilize the temperature of OpLev light source.
I asked Tanaka-kun to run the script after his work, so I will check the result in the next morning.
Tamaki-kun ran the script and started the search now.
Since MN OpLev is so glithy today, I changed the temperature from23.6 to 23.9 degrees.
Today, So many glithches occurres in EY MN TILT oplev again. so we gave up the today's IFO work...